Tim is 5 years older than Jean.Their ages in years added up to 25.How old is Tim and Jean?
1
Expert's answer
2021-10-21T06:51:48-0400

Tim's age=X,

Jean's age=Y.

Tim is 5 years older than Jean,

"\therefore" X=Y+5...eqtn 1

Their ages add up to 25,

"\therefore" X+Y=25...eqtn 2

Substitute eqtn 1 into eqtn 2

(Y+5)+Y=25

2Y+5=25

2Y=25-5

2Y=20

Y=10.

Substitute Y=10 into eqtn 1

X=Y+5

X=10+5

X=15.

"\therefore" Tim's age is 15 and Jean's age is 10.